Help would be much appreciated.
Write an equation for interest using the given data below. Use x for the unknown.
I = $95.00, P = ?, r = 5%, t = 4 years

Answer by ewatrrr(24785)

need to clarify: simple or compounded interest
I = $95.00, P = ?, r = 5%, t = 4 years
.05*P*4yr = $95
P = $95/(.05*4) = $475
CHECKING our answer***
$475 * .05*4 = $95
Wish You the Best in your Studies.